Air-saturated water has a hydronium ion concentration caused by the dissolved CO2 of 2.0 × 10–6 M, about 20-times larger than that of pure water. Calculate the pH of the solution at 25 °C.


1
Expert's answer
2021-06-07T05:03:30-0400

Concentration of the hydronium ions caused by the dissolved CO2=2 .0×10-6M

Concentration of hydronium in pure water = 2.0×10-6"\div" 20

=1.0×10-7M


Therefore, Total concentration of hydronium ions=(2.0×10-6+1.0×10-7)M

=1.0×10-7(20+1)M

=21×10-7M

=2.1×10-6M

"\therefore" pH=-log[H3O+]=-log(2.1×10-6)

=5.68

pH of the solution at 25 °C is 5.68